STUDY ON IVR OF THE N-H STRETCHING VIBRATION OF ANILINE BY PICOSECOND IR-UV PUMP-PROBE SPECTROSCOPY IN A SUPERSONIC BEAM

Y. YAMADA, N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578; \underline{T. EBATA}, Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi Hiroshima 739-8526, Japan.